What is the mass of each ore in atomic mass units?
Hematite Fe2O3

pyrite FeS2
1
Expert's answer
2011-05-18T07:35:33-0400
Fe2O3: 2*M(Fe) + 3*M(O) = 2* 56 a.m.u. + 3* 16 a.m.u. = 160 a.m.u.
FeS2: M(Fe) + 2*M(S) = 56 a.m.u. + 2* 32 a.m.u. = 120 a.m.u.